OpenStudy (anonymous):

Solve for x: 4/3x +2/3 = 6 x = 4 x = 3 x = 1 x = 0

OpenStudy (anonymous):

(4/3)x +(2/3)=6/1 (6/1)*3=18/3 (4/3)=(18/3)-(2/3)=16/3 The rest you can easily find yourself.

OpenStudy (anonymous):

@Sjano I'm Confused!

OpenStudy (anonymous):

Did you understand what i did? (Also i forgot the x at the (4/3) at the bottom)

OpenStudy (anonymous):

Isolate the x.\[\frac{4x}{3} +\frac{2}{3} = 6 \implies \frac{4x}{3} = 6-\frac{2}{3}=\frac{16}{3}\implies x=4\]